# MeanCache: User-Centric Semantic Cache for Large Language Model Based Web Services

### v3

@@ 5.4. MeanCache Comparison with Baseline
Initially, we send a new set of one thousand queries to Llama 2 (i.e., without any semantic cache) to establish a baseline for response times. We limit responses to 50 tokens to reflect practical response sizes, although actual sizes can be much larger. Note that MeanCache’s performance is not dependent on the response as it only matches the queries. Next, we send these queries to Llama 2 based local LLM service with MeanCache and GPTCache to measure the response times and performance metrics (e.g., precision, recall, F-score) respectively. An analysis of a random subset of 100 queries (70 non-duplicate and 30 duplicate) from the 1000 queries shows the impact of caching on response times in Figure 9 and cache hit and miss rates in Figure 10. The X-axis represents the query and the Y-axis represents each semantic cache hit and miss alongside the real label in Figure 10. The Y-axis in Figure 9 shows the response time of each query without any cache (Llama 2), with GPTCache (Llama 2+ GPTCache), and with MeanCache (Llama 2+MeanCache). Figure 9 demonstrates that implementing a semantic cache does not impede the performance (queries ranging from 0 to 69) and improves the user experience as response times reduces on duplicate queries (queries 70 to 99).
@@
However, Figure 10 shows that GPTCache produces significant false hits on non-duplicate queries (queries 0 to 69) compared to MeanCache. Each false hit means the user receives an incorrect query response from the cache and must resend the same query to LLM service, resulting in a poor user experience. To prioritize precision, we adjust the beta value in the F score to 0.5, valuing precision twice as highly as recall to ensure user satisfaction by avoiding false positives. This decision is driven by the need to minimize user inconvenience caused by incorrect cache hits. A false miss (low recall) does not require user intervention as the false miss query will be automatically routed to the LLM. Thus, precision in semantic caching is more important than recall.
@@
Table 1 and the confusion matrices in Figure 11 highlight MeanCache’s superior performance over GPTCache on the 1000 user queries. Notably, MeanCache with MPNet achieves a precision of 0.72, significantly surpassing GPTCache’s 0.52. This superiority is evident in the lower false positive rates (i.e., false hits) shown in Figure 11(a). The number of false hits for MeanCache is 89 (Figure 11(a)), while GPTCache has 233 false hits, as depicted in Figure 11(b). In practical terms, this means that with GPTCache, the end user has to manually resend 233 queries to the LLM service to get the correct responses, compared to only 89 queries with MeanCache. While GPTCache’s recall is higher than MeanCache’s, as we discussed earlier, precision is significantly more important than recall, and MeanCache outperforms GPTCache in this regard. Overall, the F-score of MeanCache with MPNet is 0.73 and 0.68 with the Albert embedding model, both of which outperform GPTCache’s F-score of 0.56.
@@
Summary. In an end-to-end deployment with Llama 2 as LLM service, MeanCache significantly outperforms GPTCache. It demonstrates a 17% higher F score and a 20% increase in precision in optimal configuration. The substantial reduction in false cache hits enhances the end-user experience.
5.5. Contextual Queries
@@
Figure 13. Performance on Contextual Queries: MeanCache vs. Baseline. (a) reports MeanCache’s fewer false hits 3 vs. 54 of GPTCache. (b) reports higher true hits by MeanCache.
@@
Figure 14. MeanCache only reports three false hits compared to 54 false hits by GPTCache. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
The end-user interaction with an LLM service mainly involves two types of user queries: standalone (e.g., Q1 Draw a line in Python?) and follow-up or contextual queries (e.g., Q2 Change the color to red). While standalone queries can be processed independently, contextual queries require prior context for accurate responses. Suppose that both Q1 and Q2 with their responses are cached. The user sends a new query, Q3 Draw a circle?, followed by Q4 Change the color to red. Q4 is semantically similar to Q2 in the cache but refers to a different context. GPTCache incorrectly identifies this as a cache hit, leading to an inaccurate response. MeanCache addresses this issue by verifying the context chain (Algorithm 1) of semantically matched queries. This approach ensures that contextual queries (Q4) result in a true cache miss, prompting a request to the LLM service for an appropriate response.
@@
On the dataset of 450 contextual queries (see Section 5.1.1), first, we populate the cache (MeanCache and baseline) with 200 queries (100 standalone and 100 contextual queries of the standalone queries). Next, we send 150 duplicate queries (75 standalone + 75 contextual) and 100 non-duplicate queries (a total of 250 queries) to the cache-enabled LLM. Figure 13 shows the true label (whether the query should be returned from the cache or not) and the corresponding GPTCache and MeanCache performance (predicted label). Note that in Figure 13(a), all the queries should be answered by the LLM; in other words, there should be no hits. However, GPTCache has 54 false hits, while MeanCache has only three false hits. This is also shown in the confusion matrix in Figure 14. Table 1 (Column-3) summarizes the comparative results. MeanCache outperforms GPTCache by over 25% in both F-Score and accuracy. Additionally, MeanCache achieves 32% higher precision compared to the baseline.
@@
Summary. GPTCache’s low performance stems from its inability to consider contextual information, leading to high false hit rates. In contrast, MeanCache demonstrates significantly superior handling of contextual queries, with a 25% improvement in accuracy over the baseline.
5.6. MeanCache Embedding Compression utility and Impact on Storage Space
@@
Clients or users often face storage limitations compared to web servers. Storing embeddings in the local cache on the user side for semantic search demands memory storage. Various models yield embeddings with differing vector sizes; for example, the MPNet and Albert models produce an output embedding vector of 768 dimensions, whereas the Llama 2 model’s embeddings dimension size is 4096. The embedding vector size also affects semantic search speeds, where smaller vectors could enhance speed and lower resource demands.
@@
Figure 12 illustrates the effects of MeanCache dimension reduction utility on the storage, semantic matching speed (overhead), and MeanCache’s performance (F score). The x-axis indicates the number of queries stored in the cache, while the y-axis shows storage size, average search time, and the F score in respective graphs. MeanCache-Compressed (MPNet) and MeanCache-Compressed (Albert) represent instances where MeanCache decreases the embedding dimensions from 768 to 64 by employing the compression, as detailed in MeanCache design (Section 4).
Figure 12 demonstrates that an increase in stored queries linearly raises storage needs. Yet MeanCache with compression enabled drastically lowers storage needs by 83% compared to GPTCache.
@@
Figure 12 also indicates that compression decreases the average search time, with MeanCache enabled compression approximately 11% faster. Moreover, despite a slight decrease in F score with compression enabled, MeanCache still surpasses GPTCache. Furthermore, given the evidence from Figures 5 and 6 that MPNet produces more precise embeddings, and it is also clear from Figure 12 that MPNet’s embeddings are particularly resilient to compression and excel in semantic matching.
@@
Figure 15. Llama 2 takes significantly longer to compute embeddings and requires substantially more storage space than Albert and MPNet.
@@
Figure 16. Llama 2 as an embedding model does not perform well in semantic matching, even at optimal thresholds. The maximum F1 score is 0.75, significantly lower than with the smaller embedding models (MPNet and Albert).
5.7. Infeasibility of Embedding Generation with Llama 2
GPTCache(Bang, 2023) recommend using Llama 2 for generating embeddings to enhance GPTCache’s performance. However, Llama 2’s embedding computation is expensive in terms of inference time, requires substantial storage, and incurs considerable overhead during semantic searches. For example, Llama 2 with its 7 billion parameters, demands 30 GB of memory (Siz, [n. d.]), whereas Albert and MPNet require only 43 MB and 420 MB, respectively.
@@ 5.7. Infeasibility of Embedding Generation with Llama 2
Furthermore, we evaluate the performance of the Llama 2 on embedding generation and semantic matching. Figure 16 shows that the performance of Llama 2 with different cosine similarity threshold ($\tau$) and corresponding performance metric. We can see that the performance of Llama 2 is not good even with the optimal cosine similarity threshold, as also noted by researchers (Use, [n. d.]). The maximum F1 score achieved by Llama 2 is 0.75 which is quite low when compared with the optimal thresholds scores from the Figures 7 and Figure 8.
@@
Smaller models tailored for specific tasks often surpass larger models in efficiency (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). Thus, diverging from GPTCache’s approach, we advocate for adopting smaller yet efficient embedding models for semantic caching. These models not only ensure optimal performance but also minimize the semantic cache’s overhead on users, featuring lower inference demands and reduced output embedding sizes, thereby facilitating deployment on edge devices.
@@
Summary. Llama 2 as an embedding model is not a viable option for generating embeddings. Future enhancements might improve its performance to generate high-quality embeddings, yet the computational demands, semantic search duration, and storage requirements are likely to remain elevated, consequently impacting the system’s overall efficacy.
6. Related Work
@@
Several caching systems are proposed to optimize the performance of search engines, databases and operating systems. Saraiva et al. (2001) suggest a two-tier dynamic caching architecture for web search engines to enhance response times in hierarchical systems. Utilizing LRU eviction at both levels, they demonstrate how the second-tier cache can significantly lower disk traffic and boost throughput.  Baeza-Yates and Saint-Jean (2003) propose a three-level index organization and  Long and Suel (2005) propose a three-tier caching. Xie and O’Hallaron (2002) examined two real search engine datasets to explore query locality, aiming to develop a caching strategy based on this concept. Their analysis centered on query frequency and distribution, assessing the feasibility of caching at various levels, such as server, proxy, and client side. Lempel and Moran (2003) proposed a novel caching technique, called Probability Driven Cache (PDC), to optimize the performance of search engines. PDC uses the probability of a query to be repeated to decide whether to cache the query or not. Fagni et al. (2006) proposes SDC (Static Dynamic Cache) that exploit the temporal and spatial locality present in the query stream to avoid processing the same query and thereby saving the computational resources. Baeza-Yates et al. (2007) explore efficient caching design for web search engines, comparing static and dynamic caching, and weighing caching query results against posting lists. Zhang et al. (2008) examines challenges in large search engines processing thousands of queries per second across vast document collections. The study focuses on index compression and caching optimizations, evaluating various inverted list compression algorithms, including novel variants, and analyzing different caching policies (e.g., LRU, LFU). Marin et al. (2010) introduces an innovative cache hierarchy aimed at boosting web search engines’ efficiency in processing user queries. It aims to store different pieces of data useful for solving frequent queries, thereby enhancing overall query throughput and reducing single-user query latency and power consumption.
@@
In all of the above studies, the caching systems are designed for traditional search engines that process keyword queries (i.e., key word matching) and return a list of links to the user as a response. However, such caching systems, when applied to LLMs-based web servers or APIs, do not provide a single concise response and may return many false results. On top of that, such caching techniques will fail to capture the semantic similarity among repeated queries and will result in a significantly low hit rate. Zhu et al. (2023) and (Bang, 2023) propose server-side caching for the services based on LLMs to reduce the massive computational cost of LLMs. For instance, Zhu et al. (2023) upon receiving a new query at the server checks if the query is semantically similar to any query in the cache. If the query is similar to any query in the cache then the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query.
@@
Although these techniques can handle the semantic similarity among the queries and can provide a single concise response, they raise privacy concerns as they store user queries on an external server. These techniques are static and unable to adapt to the behavior of each user. Furthermore, an end user will still be charged for the query even if the query is served from the cache. Thus, we need a user-centered semantic cache that can run on the user side, which can benefit the user in terms of privacy, cost, and latency. Such a cache should be able to detect the semantic similarity among the queries and should be able to adapt to the behavior of each user in a privacy preserving manner. MeanCache provides the aforementioned benefits without violating the user’s privacy.
7. Conclusion

### v4

@@ IV-B MeanCache Comparison with Baseline
Initially, we send a new set of one thousand queries to Llama 2 (i.e., without any semantic cache) to establish a baseline for response times. We limit responses to 50 tokens to reflect practical response sizes, although actual sizes can be much larger. Note that MeanCache’s performance is not dependent on the response as it only matches the queries. Next, we send these queries to Llama 2 based local LLM service with MeanCache and GPTCache to measure the response times and performance metrics (e.g., precision, recall, F-score), respectively. An analysis of a random subset of 100 queries (70 non-duplicate and 30 duplicate) from the 1000 queries shows the impact of caching on response times in Figure 5 and cache hit and miss rates in Figure 6. While MeanCache is evaluated on 1,000 standalone and 450 contextual queries (Table I), we use this smaller subset of 100 queries solely for visualization purposes, as displaying all queries would create cluttered figures. The y-axis in Figure 5 shows the response time of each query without any cache (Llama 2), with GPTCache (Llama 2 + GPTCache), and with MeanCache (Llama 2 + MeanCache). In Figure 6, x-axis represents the query and the y-axis represents each semantic cache hit and miss alongside the real label. Figure 5 demonstrates that implementing a semantic cache does not impede the performance (queries ranging from 0 to 69) and improves the user experience as response times reduce on duplicate queries (queries 70 to 99).
@@
However, Figure 6 shows that GPTCache produces significant false hits on non-duplicate queries (queries 0 to 69) compared to MeanCache. Each false hit means the user receives an incorrect cached response, requiring them to resend the query to the LLM service, leading to a poor user experience. To prioritize precision, we adjust the beta value in the F score to 0.5, valuing precision twice as highly as recall to ensure user satisfaction by avoiding false positives. This decision is driven by the need to minimize user inconvenience caused by incorrect cache hits. A false miss (low recall) does not require user intervention as the false miss query will be automatically routed to the LLM. Thus, precision in semantic caching is more important than recall.
@@
Table I and the confusion matrices in Figure 7 highlight MeanCache’s superior performance over GPTCache on the 1000 user queries. Notably, MeanCache with MPNet achieves a precision of 0.72, significantly surpassing GPTCache’s 0.52. This superiority is evident in the lower false positive rates (i.e., false hits) shown in Figure 7(a). The number of false hits for MeanCache is 89 (Figure 7(a)), while GPTCache has 233 false hits, as depicted in Figure 7(b). In practical terms, this means that with GPTCache, the end user has to manually resend 233 queries to the LLM service to get the correct responses, compared to only 89 queries with MeanCache. While GPTCache’s recall is higher than MeanCache’s, as we discussed earlier, precision is significantly more important than recall, and MeanCache outperforms GPTCache in this regard. Overall, the F-score of MeanCache with MPNet is 0.73 and 0.68 with the Albert embedding model, both of which outperform GPTCache’s F-score of 0.56.
@@
Summary. In an end-to-end deployment, MeanCache significantly outperforms GPTCache. It demonstrates a 17% higher F score and a 20% increase in precision in optimal configuration. The substantial reduction in false cache hits enhances the end-user experience.
IV-C Contextual Queries
@@
Figure 8: Performance on Contextual Queries: MeanCache vs. Baseline. (a) reports MeanCache’s fewer false hits 3 vs. 54 of GPTCache. (b) reports higher true hits by MeanCache.
@@
Figure 9: For contextual queries, MeanCache reports three false hits, compared to 54 false hits by GPTCache.
Section II describes contextual queries. GPTCache, lacking the capability to detect contextual queries, incorrectly identifies such queries as cache hits, resulting in inaccurate responses. MeanCache addresses this limitation by verifying the context chain (Algorithm 1) of semantically matched queries. This verification ensures that contextual queries (e.g., Q4 in Section II) correctly yield a true cache miss, thus prompting an appropriate request to the LLM service.
@@
On the dataset of 450 contextual queries (see Section IV-A1), first, we populate the cache (MeanCache and baseline) with 200 queries (100 standalone and 100 contextual queries of the standalone queries). Next, we send 150 duplicate queries (75 standalone + 75 contextual) and 100 non-duplicate queries (a total of 250 queries) to the cache-enabled LLM. Figure 8 shows the true label (whether the query should be returned from the cache or not) and the corresponding GPTCache and MeanCache performance (predicted label). Note that in Figure 8(a), all the queries should be answered by the LLM; in other words, there should be no hits. However, GPTCache has 54 false hits, while MeanCache has only three false hits. This is also shown in the confusion matrix in Figure 9. Table I (Column-3) summarizes the comparative results. MeanCache outperforms GPTCache by over 25% in both F-Score and accuracy. Additionally, MeanCache achieves 32% higher precision compared to the baseline.
@@
Summary. GPTCache’s low performance stems from its inability to consider contextual information, leading to high false hit rates. MeanCache demonstrates superior handling of contextual queries, with a 25% improvement in accuracy over the baseline.
Figure 10: (a) MeanCache’s embedding compression reduces storage by 83% compared to GPTCache. (b) MeanCache’s semantic matching speed is 11% faster with compression enabled, while still outperforming GPTCache. (c) MeanCache’s F score is slightly lower with compression enabled, but it still outperforms GPTCache.
IV-D Embedding Compression and Impact on Storage Space
@@
Clients often face storage limitations compared to web servers. Storing embeddings in the local cache on the user side for semantic search demands memory storage. Various models yield embeddings with differing vector sizes; for example, the MPNet and Albert models produce an output embedding vector of 768 dimensions, whereas the Llama 2 model’s embeddings dimension size is 4096. The embedding vector size also affects semantic search speeds, where smaller vectors could enhance speed and lower resource demands.
@@
Figure 10 illustrates the effects of MeanCache dimension reduction utility on the storage, semantic matching speed (overhead), and MeanCache’s performance (F score). The x-axis indicates the number of queries stored in the cache, while the y-axis shows storage size, average search time, and the F score in respective graphs. MeanCache-Compressed (MPNet) and MeanCache-Compressed (Albert) represent instances where MeanCache decreases the embedding dimensions from 768 to 64 by employing the compression, as detailed in MeanCache design (Section III).
@@
Figure 10 demonstrates that increased stored queries linearly raise storage needs. Yet MeanCache with compression enabled drastically lowers storage needs by 83% compared to GPTCache. Figure 10 also indicates that compression decreases the average search time, with MeanCache enabled compression approximately 11% faster. Moreover, despite a slight decrease in F score with compression enabled, MeanCache still surpasses GPTCache. Furthermore, given the evidence from Section IV-E (Figures 11 and 12) that MPNet produces more precise embeddings, and it is also clear from Figure 10 that MPNet’s embeddings are particularly resilient to compression and excel in semantic matching.
@@
Figure 11: MPNet’s FL training helps generate high-quality embeddings.
Figure 12: FL training boosts MeanCache’s query matching precision with Albert.
IV-E Privacy Preserving Embeddings Model Training
Storing clients queries on the server side presents a potential privacy risk. To address this, each client can retain its local data on its own device. The ensuing challenge is how to train an embedding model that can also utilize the distributed data from all clients. FL is recognized for training neural networks in a privacy-preserving manner. As such, MeanCache employs FL to train and fine-tune an embedding model, thereby preserving privacy and leveraging the dataset residing on the client’s side. In this section, our objective is to evaluate whether FL training can progressively enhance the embedding model to generate high-quality embeddings for user queries. To simulate this scenario, we distribute the training dataset among 20 clients. In each round, we sample 4 clients, conducting a total of 50 FL training rounds. Each client trains its embedding model for 6 epochs, operating on a dedicated A100 GPU. We conduct two experiments with the Albert and MPNet models. The batch size is set to 256 for the Albert model, and for MPNet, it is set to 128 during the local training by participating clients.
Figures 11 and 12 depict the performance of MeanCache as FL training progresses. The x-axis represents the training round, while the y-axis shows the performance metrics such as F-score, precision, recall, and accuracy of the global model ($W_{global}^{t+1}$). As illustrated in Figure 11, the F-score for MPNet increases from 0.82 to 0.88, and for Albert, it rises from 0.83 to 0.86, as shown in Figure 12. Similarly, precision for MPNet significantly increases from 0.74 to 0.85, as depicted in Figure 11, and for Albert, it increases from 0.74 to 0.81, as demonstrated in Figure 12. Given that MPNet is a more robust transformer architecture compared to Albert, it is also observed during our training that MPNet outperforms Albert, exhibiting superior learning in FL settings.
Summary. FL training increases 11% precision of MeanCache for MPNet and a 7% increase for Albert. The performance of the embedding model to generate high-quality embeddings can improve in a privacy-preserving manner using FL training.
Figure 13: MeanCache optimizes MPNet’s threshold.
Figure 14: MeanCache identifies an optimal threshold of 0.78 for Albert.
IV-F Cosine Similarity Threshold Impact on Semantic Matching
Semantic matching for a new user query begins by generating the embeddings of the user’s query ($E_{q}$) using the embedding model. The cosine similarity ($\theta$) is computed with the cached embeddings. If the cosine similarity $\theta$ exceeds the threshold $\tau$, the cache is hit and the response to the user query is returned from the cache. Therefore, the cosine similarity threshold $\tau$ is crucial in determining the similarity between a user query and cached entries. A low threshold value of $\tau$ can lead to false hits (incorrect matches), while a high threshold might overlook the appropriate matches (i.e., false cache misses or false negatives).
To illustrate this, MeanCache varies the threshold $\tau$ from 0 to 9 and evaluates the performance metrics F-score, precision, recall, and accuracy with an equal distribution of duplicate and non-duplicate queries from the validation data to avoid bias. Figures 13 and 14 show how the cosine similarity threshold ($\tau$) affects MeanCache’s performance. The x-axis represents the threshold $\tau$ values, and the y-axis denotes the performance metrics. For instance, at a 0.3 threshold, MeanCache’s semantic matching accuracy using MPNet is 57%, with a precision of 54% as shown in Figure 13. Similarly, with Albert at the same threshold, the accuracy is 55%, and the precision is 53% (Figure 14). Precision typically improves with an increase in threshold. However, beyond a certain point, increasing the threshold $\tau$ leads to a decline in F score, accuracy, and recall.
For MPNet, the optimal threshold $\tau$ is identified at 0.83, achieving an F1 score of 0.89, precision of 0.92, and accuracy of 0.90 (Figure 13). For Albert, the optimal threshold is 0.78, with an F1 score of 0.88.
Summary. GPTCache’s suggested threshold of 0.7 will result in suboptimal performance during semantic matching. The optimal threshold $\tau$ values varies with the embedding model. MeanCache optimally adjusts the threshold based on user data, outperforming GPTCache’s suggested threshold by 16% in precision and 4% in F score for MPNet, and by 10% in precision and 2% in F score for Albert.
IV-G Infeasibility of Embedding Generation with Llama 2
GPTCache [20] recommend using Llama for generating embeddings to enhance GPTCache’s performance. However, Llama 2’s embedding computation is expensive in terms of inference time, requires substantial storage, and incurs considerable overhead during semantic searches. For example, Llama 2 with its 7 billion parameters, demands 30 GB of memory [50], whereas Albert and MPNet require only 43 MB and 420 MB, respectively.
@@ IV-G Infeasibility of Embedding Generation with Llama 2
Figure 15: Llama 2 takes significantly longer to compute embeddings and requires substantially more storage space than Albert and MPNet.
@@
Figure 16: Llama 2 does not perform well in semantic matching, even at optimal thresholds.
@@
Furthermore, we evaluate the performance of the Llama 2 on embedding generation and semantic matching. Figure 16 shows that the performance of Llama 2 with different cosine similarity threshold ($\tau$) and corresponding performance metric. We can see that the performance of Llama 2 is not good even with the optimal cosine similarity threshold, as also noted by researchers [51]. The maximum F1 score achieved by Llama 2 is 0.75 which is quite low when compared with the optimal thresholds scores from the Figures 13 and Figure 14.
@@
Smaller models tailored for specific tasks often surpass larger models in efficiency [23, 24, 25]. Thus, diverging from GPTCache’s approach, we advocate for adopting smaller yet efficient embedding models for semantic caching. These models not only ensure optimal performance but also minimize the semantic cache’s overhead on users, featuring lower inference demands and reduced output embedding sizes, thereby facilitating deployment on edge devices.
@@
Summary. Llama 2 is not viable for generating embeddings. Future enhancements might improve its performance to generate embeddings, but the computational demands, semantic search duration, and storage requirements will likely remain elevated.
V Related Work
@@
Several caching systems are proposed to optimize the performance. Study [18] suggests a two-tier dynamic caching architecture for web search engines to enhance response times in hierarchical systems. Utilizing LRU eviction at both levels, they demonstrate how the second-tier cache can significantly lower disk traffic and boost throughput. Researchers in [52] propose a three-level index organization and [53] propose a three-tier caching. Another study [14] examined two real search engine datasets to explore query locality, aiming to develop a caching strategy based on this concept. Their analysis centered on query frequency and distribution, assessing the feasibility of caching at various levels, such as server, proxy, and client side. A novel caching technique called Probability Driven Cache (PDC) is proposed in [13] to optimize the performance of search engines by using the probability of query repetition to decide whether to cache the query. PDC uses the probability of a query to be repeated to decide whether to cache the query or not. A different approach is presented in [54], which proposes the Static Dynamic Cache (SDC) to exploit temporal and spatial locality present in the query stream, avoiding redundant processing and saving computational resources. Efficient caching designs for web search engines are explored in [17], where static and dynamic caching strategies are compared, weighing the benefits of caching query results against posting lists. Challenges in large-scale search engines, which process thousands of queries per second across vast document collections, are examined in [55], focusing on index compression, caching optimizations, and evaluating various inverted list compression algorithms alongside caching policies such as LRU and LFU.
@@
All of these studies focus on caching systems designed for traditional search engines that process keyword queries (i.e., keyword matching) and return a list of links as a response. However, when applied to LLM-based web servers or APIs, these caching systems do not provide a single concise response and may yield many false results. Moreover, such caching techniques fail to capture the semantic similarity among repeated queries, leading to a significantly low hit rate. Server-side caching for services based on LLMs is proposed in [19] and [20], aiming to reduce the massive computational cost of LLMs. In particular, the approach in [19] checks if a new query is semantically similar to any existing queries in the cache. If a match is found, the cached response is returned; otherwise, a model multiplexer selects the most suitable LLM.
@@
While these techniques can handle semantic similarity among queries and provide a single concise response, they raise privacy concerns as user queries are stored on external servers. Additionally, these techniques are static and unable to adapt to individual user behavior. Users may still be charged for the query, even if it is served from the cache. Therefore, a user-centered semantic cache that operates on the user side is needed, providing benefits in terms of privacy, cost, and latency. This cache should be able to detect semantic similarity among queries and adapt to each user’s behavior while preserving privacy. MeanCache offers these benefits without compromising user privacy.
VI Conclusion